OpenStudy (anonymous):

if P is prime then for what values of P expression 4P+1 yeilds a perfect square?

OpenStudy (anonymous):

the number whose square it is, must be an odd number \[4P=a^2-1\\ P=(a/2)^2-(1/4)\] 1. "a" is odd 2. P = even ? only for P=2?

OpenStudy (anonymous):

\[P=\left({a\over2}+{1\over2}\right)\left({a\over2}-{1\over2}\right)\] P is a mean of two odds. which makes it even

OpenStudy (anonymous):

@wio please verify

OpenStudy (anonymous):

Thats Good. Answer is 2..Thanks

OpenStudy (anonymous):

That looks correct. You could have kept the 4 factored out though.

OpenStudy (anonymous):

Though I'm not sure what they want when they say 'what values of P'

OpenStudy (anonymous):

@hashim12atd how about some appreciation with a medal then?

OpenStudy (anonymous):

@wio know.. but the steps tell me that P must be even!

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

ok so we are ok till "a=odd"?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

\[P=(a/2+1/2)(a/2−1/2) \] one of these factors is inevitably even . So P is even. or P=2. a= even gets us no where, as it doesn't yield an integer

OpenStudy (anonymous):

He said \(a\) is odd. Replace \(a\) with \(2n+1\)

OpenStudy (anonymous):

so, from the third relation, we conclude, (odd/2+0.5) = some number and (odd/2-0.5) = another number should always give us an even

OpenStudy (anonymous):

@hashim12atd and @wio me and my twisted schemas..

OpenStudy (anonymous):

@electrokid Yeah whenever you claim something is even or odd, then use the \(2n\) and\(2n+1\) paradigm to sell it. I knew before hand even times even is even and odd times odd is odd, but that isn't always obvious.

OpenStudy (anonymous):

how we award medals?

OpenStudy (anonymous):

click best response.

OpenStudy (anonymous):

@wio i know, i was trying a couple of ways and doing things mentally and lost track of penning them down

OpenStudy (anonymous):

okay man..

OpenStudy (anonymous):

where from are you? electrokid

OpenStudy (anonymous):

Pakistan.